The perimeter of a rectangle is 40.
If the width is 3x-8 and the length 2x + 3 what is x?

Answer by checkley77(12844)

W=3x-8
L=2x + 3
2L+2W=40
2(2X+3)+2(3X-8)=40
4X+6+6X-16=40
10X=40-6+16
10X=50
X=50/10
X=5 SOLUTION.
W=3*5-8=15-8=7 ANS.
L=2*5+3=10+3=13 ANS.
PROOF:
2*13+2*7=40
26+14=40
40=40